# Self-Specialization: Uncovering Latent Expertise within Large Language Models

### v1

@@ Abstract
Recent works have demonstrated the effectiveness of self-alignment in which a large language model is, by itself, aligned to follow general instructions through the automatic generation of instructional data using a handful of human-written seeds. Instead of general alignment, in this work, we focus on self-alignment for expert domain specialization (e.g., biomedicine), discovering it to be very effective for improving zero-shot and few-shot performance in target domains of interest. As a preliminary, we first present the benchmark results of existing aligned models within a specialized domain, which reveals the marginal effect that “generic” instruction-following training has on downstream expert domains’ performance. To remedy this, we explore self-specialization that leverages domain-specific unlabelled data and a few labeled seeds for the self-alignment process. When augmented with retrieval to reduce hallucination and enhance concurrency of the alignment, self-specialization offers an effective (and efficient) way of “carving out” an expert model out of a “generalist”, pre-trained LLM where different domains of expertise are originally combined in a form of “superposition”. Our experimental results on a biomedical domain show that our self-specialized model (30B) outperforms its base model, MPT-30B by a large margin and even surpasses larger popular models based on LLaMA-65B, highlighting its potential and practicality for specialization, especially considering its efficiency in terms of data and parameters.
@@ 1 Introduction
Instruction-tuning (Ouyang et al., 2022; Wei et al., 2022; Mishra et al., 2022; Su et al., 2022) of large language models (LLMs) offers a mechanism to adeptly guide models using specific directives, thereby enhancing their versatility across diverse tasks. However, as promising as this concept might seem, it poses an inherent challenge: the substantial need for quality data (Chung et al., 2022; Wan et al., 2023; Köpf et al., 2023). The very premise of instruction-tuning hinges on the availability of well-crafted, human-annotated data, a resource that is both time-consuming and challenging to scale efficiently (Honovich et al., 2022; Kang et al., 2023). Furthermore, acquiring domain-specific data is even more demanding as it requires the involvement of domain experts, which is often more expensive (Bai et al., 2021; Wang et al., 2023).
@@
Emerging as a promising solution to this data-intensive challenge is the approach of self-alignment (Wang et al., 2022a; Sun et al., 2023). By allowing LLMs to automatically generate instructional data from a handful of human-authored seeds, self-alignment presents a means to harness the internal general knowledge of these models (which results from extensive pre-training on the internet corpora (Devlin et al., 2019; Raffel et al., 2020; Brown et al., 2020)) without extensive human annotations.
@@
However, some pertinent questions remain: (i) How effective are the original or the self-aligned models when applied to more niche domains, such as biomedicine? (ii) Given that neither the initial pre-training nor subsequent self-alignment tuning is domain-specific, we hypothesize that the model expertise in different domains resides in “superposition” in the model’s parameters and hidden states. In other words, parametric knowledge in LLMs represents a mixture of semantics and knowledge of various domains. May this hinder each individual expert-domain performance? These inquiries become even more practical when considering the ever-growing demands for specialized models that can cater to domain-specific nuances. In our preliminary study, we find that existing models such as Alpaca (Taori et al., 2023) and Dromedary (Sun et al., 2023), although aligned, exhibit only a modest degree of improvement (compared to source models before alignment) within the specialized domains. These observations underline the need for innovative approaches that can harness the depth of domain-specific knowledge existing in the base models, to ensure the self-generated instruction-tuning data remains both contextually appropriate and accurate.
@@
In this work, we introduce and explore the new concept of self-specialization (Fig. 1). Drawing inspiration from the foundational principles of self-alignment, self-specialization goes a step further by incorporating domain-specific seeds and external knowledge. Our approach integrates specialized seed instructions and is further bolstered by a retrieval component. Our goal is to guide models beyond generic alignment, directing them to generate data that are not just contextually fitting for a specialized domain but also maintain high degrees of accuracy.
Through rigorous experiments, we evaluate our self-specialized models within the biomedical domain. Surprisingly, despite the apparent simplicity of our approach, our results, present a compelling case for self-specialization. The experimental results demonstrate that our self-specialized model (30B) using this approach outperforms its base model, MPT-30B (Team, 2023) by a large margin and even surpasses larger models (based on LLaMA-65B (Touvron et al., 2023)), including the ones improved through self-alignment by leading methods (Taori et al., 2023; Sun et al., 2023). Moreover, we show that effective self-alignment is possible with parameter-efficient finetuning techniques (PEFT (Mangrulkar et al., 2022), QLoRA (Dettmers et al., 2023) in our case). This opens an exciting opportunity for memory-efficient multi-specialized models, where we envision that the base LLM is loaded once and is “surrounded” by a set of low-memory-footprint LoRA modules, delivering all the specialization of interest (and potentially their mix (Huang et al., 2023)) at once and in-memory without re-loading (Fig. 1).
Consequently, the contributions of our work encompass:
We conduct comprehensive benchmarking of general-purpose aligned models within a specialized domain, underscoring the intrinsic challenge of encoding vast general knowledge into a finite set of parameters, motivating the need for specialization.
This work explores a lightweight solution, self-specialization that enables us to uncover latent expertise within LLMs with minimal supervision.
Our experiments in a biomedical domain demonstrate the remarkable potential of self-specialization, showcasing its efficiency and practicality. The promising results, achieved with this simple scheme, open new avenues for future work in this realm.
@@ 2 Preliminaries: Benchmarking Existing Aligned Models
To motivate our exploration of self-specialization, we first begin by addressing a fundamental question: How well do generally aligned models perform on specialized domains? While existing popular ones such as Alpaca and Dromedary have demonstrated the generalizability of following instructions in a general scenario, it remains unclear whether general alignment can also elicit expertise for a certain domain.
Investigating this, we assess the capabilities of Alpaca (Taori et al., 2023) and Dromedary (Sun et al., 2023) against a base model, LLaMA (Touvron et al., 2023), on a collection of benchmarks within the biomedical domain. Ensuring an unbiased comparison, all models are equally fixed with 65B parameters and share the same architecture (LLaMA). We select 10 different biomedical NLP datasets, covering a diverse set of tasks to ensure a comprehensive mix of content and also to look at the cross-task generalization, the core of instruction tuning. Few-shot (k=5) settings are examined where demonstrations are tailored to each task. Note that in fact, Alpaca is not “self”-aligned in that it uses datasets generated by GPT-3.5 (Ouyang et al., 2022) following the self-instruct process (Wang et al., 2022a), unlike Dromedary which uses the same base model. Nonetheless, we also benchmark it to serve a sort of upper bound. The results are shown in Figure 2. Details of the setups are described in Section 4.1.
When benchmarked on the expert (biomedical) domain, we find that both Alpaca and Dromedary have only a slight (1.1 - 2.5%) advantage over LLaMA. While they are aligned to handle a broad set of instructions, they do not seem to effectively improve their specialized domain expertise; intuitively trading their expertise for generality given finite parameters. In light of these findings, it becomes evident that for cases where we are only interested in target domains for all our downstream tasks, there exists a substantial potential for enhancement within domain-specific tasks. This underscores the need for a model or approach, like self-specialization, that could potentially uncover specialization while maintaining cross-task generalizability with minimal supervision. Moreover, if the approach is parameter efficient, it constitutes a considerable advantage, as many specialized models can be efficiently served in memory, sharing their deployment on the same machines / GPUs.
@@ 3 Self-Specialization
In this section, we describe our method called self-specialization illustrated in Figure 3. Starting with a select set of human-crafted, domain-specific seed instructions, the base model evolves to generate synthetic instructions and corresponding input contexts intrinsic to the domain. As we progress to the response generation phase, we enhance responses with domain-centric knowledge accessed via a retrieval mechanism, retrieving instruction-related knowledge from a domain-specific, and yet unlabeled, source. Upon generation, the specialization triggering stage follows where the base model is aligned, calibrating its expertise to resonate with the target domain. Optionally, this process can be reiterated with the aligned model as a better generator.
3.1 Seed instructions
At the outset, we harness a curated set of seed instructions , consisting of a triplet , comprised of instruction , a context (e.g., passage), and a response , respectively. Considering the real-world scenarios where domain-specific data are relatively harder to acquire (Bai et al., 2021), we aim to have a very minimal number of seed instructions. For example, we use only 80 seeds for the biomedical domain, where we sample from existing biomedical NLP benchmarks (Parmar et al., 2022) (detailed in Section 4.1). These seeds encapsulate the “spirit” of the fundamental concepts and intricacies of the targeted domain and yet are clearly insufficient to encompass the entirety of domain knowledge. We conjecture (and surprisingly demonstrate), that the domain knowledge is already residing inside the sufficiently large models generally pre-trained, yet is in a state of superposition, which does not however prevent this knowledge from being uncovered, e.g. by the means of the proposed approach. The instruction seeds are pivotal, acting as the launching pad for the model’s journey into specialization.
@@ 3.2 Domain-Specific Instruction Generation
With the seed instructions in place, we move to generating domain-specific instructions. While these new instructions are grounded in the initial seeds, they grow to cover a comprehensive scope of the domain. Specifically, a base model , such as MPT-30B (Team, 2023) which is “large enough”, is prompted to produce new combinations of given a handful of seed demonstrations which are randomly sampled from the initial seeds pool. The newly formed instructions , coupled with their corresponding input contexts , shape a blueprint that the model utilizes in the following stages.
@@ 3.3 Domain-Specific Response Generation
Once the synthetic domain-specific instructions and their corresponding contexts are in hand, our approach navigates to the response generation phase. It is certainly imperative for the response not only to be correct but also to be well-aligned with the target domain. We posit that incorporating external domain-relevant knowledge would be beneficial for this case, inspired by Frisoni et al. (2022). Therefore, we let the model also leverage a retrieval component, to infuse its responses with external, domain-relevant knowledge retrieved from an unlabeled domain-specific collection of documents. Specifically, forming the query as a concatenation of and , a retriever fetches top- relevant documents .
Then, each document is independently paired with the query to form a prompt to , and the final domain-specific responses are produced from the final distribution computed by marginalizing over the probabilities of each of these -combinations at each generation step. The objective of this generation can be thus formally represented as:
where is a relevance score (similarity) from a retriever module. By integrating such domain-specific information, this step encourages the generated target responses to be more nuanced and domain-specific leading to improvements (Section 4.3).
@@ 3.4 Triggering Specialization
Upon establishing a robust set of domain-specific responses, the model enters the specialization phase. Here, it undergoes tuning using the synthetic instructional data generated by itself, adjusting its internal parameters (i.e., QLoRA (Dettmers et al., 2023)) to cater specifically to the domain’s nuances. This step is crucial, marking the model’s transformation from being generally competent to being domain-specialized while preserving cross-task generalizability, thus resulting in the final self-aligned domain-specialized model: .
@@ 3.5 Iterative Self-Specialization
In the spirit of continuous improvement, our approach optionally undergoes iterative self-specialization. It revisits the generation process of instructions and responses with the better-aligned model . Here, to maximize the effectiveness of this process, we re-visit the idea of a contrastive decoding scheme (Li et al., 2023). The original idea of Li et al. (2023) is that a larger model can be contrasted with a smaller model in output distributions to improve the generation quality of the larger model. In our case, is adopted for the stronger model, whereas is considered as the weaker model, which can be represented as:
This process has the potential of refining the model’s domain expertise with each iteration (of considering the previous iteration as base each time), iteratively improving its responses.
4 Experiments
4.1 Experimental Setups
@@
For evaluation, we employ various biomedical NLP datasets, most of which are curated in BigBio (Fries et al., 2022). A total of 10 different datasets are adopted to encompass a wide range of NLP tasks: Question Answering (QA), Named Entity Recognition (NER), Relation Extraction (RE), Sentiment Analysis (SA), and Document Classification (DC). Following a prior work (Parmar et al., 2022), all datasets are transformed into instructional data. Specifically, we employ datasets including BioASQ-8b (Factoid, List, Yesno) (Nentidis et al., 2020), PubMedQA-Long (Jin et al., 2019), AnatEM (Pyysalo & Ananiadou, 2013), BioNLP13CG (Pyysalo et al., 2013), NCBI (Dogan et al., 2014), DDI (Herrero-Zazo et al., 2013), Medical Drugs (Khan, 2019), and HoC (Baker et al., 2015). Further details on each of these datasets are elaborated in Appendix A.
@@
We use base MPT model (Team, 2023), one of the powerful open-source foundation models, especially due to its feature of 8k context length. Inspired by the success of a previous work (Sun et al., 2023) that showed that large model size has a significant effect, we specifically adopt the 30B variant for our main experiments. For the retriever, we use simple yet effective BM25 (Robertson et al., 1994), in order to support a practical scenario where sufficient human-labeled data for training a more sophisticated retriever is not available. In addition to MPT-30B, we evaluate Falcon-40B (Almazrouei et al., 2023), another strong open-source model, to further validate the general applicability of self-specialization. For benchmarking of general-purpose aligned models, we evaluate Alpaca-65B (Taori et al., 2023) and Dromedary-65B (Sun et al., 2023) that are both built based on LLaMA (Touvron et al., 2023).
@@
For domain-specific seed data, we use data sampled from BoX (Parmar et al., 2022), which encompasses 32 tasks, up to 5 instances for each dataset, resulting in a compact yet representative seed data of 80 samples in total. These seeds are also used as demonstrations in a prompt for inference. For domain-specific external corpus, we leverage PubMed111gs://scifive/pretrain/pubmed_cleaned preprocessed in (Phan et al., 2021), which contains 30M abstracts. We generate 5K synthetic instructional data through the self-specialization process. Being equipped with QLoRA (Dettmers et al., 2023) and 4-bit quantization, the model is trained using a simple Alpaca-style template (Taori et al., 2023) on a single A100, taking only a few hours for 3 epochs, resulting in a light-weight (parameter efficient) specialization module that can be attached to the base model inducing its specialization upon request.
-Score k=0 k=1 k=5 Task Dataset Base Self-Specialized Base Self-Specialized Base Self-Specialized QA BioASQ-Factoid 30.90 37.35 47.56 55.04 51.96 57.61 BioASQ-List 46.06 46.99 47.57 44.55 35.09 42.17 BioASQ-Yesno 21.20 85.27 10.80 94.00 8.80 95.20 PubMedQA 11.98 24.16 28.89 24.87 31.69 31.31 Average 27.54 48.44 33.71 54.62 31.89 56.57 NER AnatEM 9.63 11.99 7.57 15.76 6.59 21.25 BioNLP13CG 24.79 24.93 21.76 31.80 26.03 41.16 NCBI 18.46 14.35 27.88 43.11 17.99 46.54 Average 17.63 17.09 19.07 30.22 16.87 36.32 RE DDI 51.00 49.40 49.20 51.60 49.38 53.40 SA Medical Drugs 35.00 65.80 11.40 54.60 11.40 32.80 DC HoC 2.44 6.01 13.91 7.61 62.84 62.65 Average 25.15 36.63 26.65 42.29 30.18 48.41
Rouge-L k=0 k=1 k=5 Task Dataset Base Self-Specialized Base Self-Specialized Base Self-Specialized QA BioASQ-Factoid 30.70 37.31 47.35 54.71 51.81 57.48 BioASQ-List 41.07 40.65 42.38 38.50 30.40 36.24 BioASQ-Yesno 21.20 85.27 10.80 94.00 8.80 95.20 PubMedQA 9.15 18.88 22.78 18.52 24.56 24.77 Average 25.53 45.53 30.83 51.43 28.89 53.42 NER AnatEM 8.65 10.69 6.67 13.83 6.07 19.24 BioNLP13CG 20.41 20.34 19.02 27.54 22.53 35.07 NCBI 17.94 13.75 25.22 39.27 16.60 41.55 Average 15.86 14.93 16.97 26.88 15.07 31.95 RE DDI 51.00 49.40 49.20 51.60 49.38 53.40 SA Medical Drugs 35.00 65.80 11.40 54.60 11.40 32.80 DC HoC 2.42 5.83 13.88 7.61 62.84 62.61 Average 23.75 34.79 24.87 40.02 28.44 45.84
4.2 Results

### v2

@@ Abstract
Recent works have demonstrated the effectiveness of self-alignment in which a large language model is aligned to follow general instructions using instructional data generated from the model itself starting from a handful of human-written seeds. Instead of general alignment, in this work, we focus on self-alignment for expert domain specialization (e.g., biomedicine, finance). As a preliminary, we quantitively show the marginal effect that generic instruction-following training has on downstream expert domains’ performance. To remedy this, we propose self-specialization - allowing for effective model specialization while achieving cross-task generalization by leveraging only a few labeled seeds. Self-specialization offers a data- and parameter-efficient way of “carving out” an expert model out of a generalist pre-trained LLM. Exploring a variety of popular open large models as a base for specialization, our experimental results in both biomedical and financial domains show that our self-specialized models outperform their base models by a large margin, and even larger models that are generally instruction-tuned or that have been adapted to the target domain by other means.
Self-Specialization:
Uncovering Latent Expertise within Large Language Models
@@ 1 Introduction
Instruction-tuning (Ouyang et al., 2022; Wei et al., 2022; Mishra et al., 2022; Su et al., 2022) of large language models (LLMs) offers a mechanism to adeptly guide models using specific directives, thereby enhancing their versatility across diverse tasks. However, as promising as this concept might seem, it poses an inherent challenge: the substantial need for quality data (Chung et al., 2022; Wan et al., 2023; Köpf et al., 2023). The very premise of instruction-tuning hinges on the availability of well-crafted, human-annotated data, a resource that is both time-consuming and challenging to scale efficiently (Honovich et al., 2022; Kang et al., 2023).
When it comes to specialized domains, such as biomedicine, it is more challenging to acquire human labels, due to the need for expert annotators (Wang et al., 2023b). While adaptation through in-domain pre-training (Gururangan et al., 2020; Wu et al., 2023) has been shown to be effective, this approach requires extensive (unlabeled) target-domain data, in addition to significant computational resources. Moreover, prior work has shown the benefits of adaptive pre-training can be less than those achieved by moderate amounts of fine-tuning data from the target domain (Bai et al., 2021).
@@
Emerging as a promising solution to this data-intensive challenge in the context of instruction-tuning is the approach of self-alignment (Wang et al., 2022a; Sun et al., 2023). By allowing LLMs to automatically generate instructional data from minimal human-authored seeds, self-alignment presents a means to harness the internal general knowledge of models, which results from extensive pre-training on internet corpora (Devlin et al., 2019; Raffel et al., 2020; Brown et al., 2020), without extensive human annotations.
@@
However, a pertinent question remains: How effective are the self-aligned models when applied to more niche domains, such as biomedicine? Given that both the initial pre-training and subsequent self-alignment are general, the knowledge embedded in LLM parameters may be a mixture of semantics and various domains. This raises questions about their effectiveness in specialized domains, despite the aims of instruction-tuning and self-alignment for cross-task generalization. In our preliminary study, however, we find that existing models such as Alpaca (Taori et al., 2023) and Dromedary (Sun et al., 2023), although aligned, exhibit only a modest degree of improvement within the specialized domains. These observations underline the need for focused approaches that can leverage the domain expertise existing in the base models, to ensure the self-generated instruction-tuning data remains both contextually appropriate and accurate.
@@
In this work, we explore the possibility of self-specialization (Fig. 1). Drawing inspiration from the foundational principles of self-alignment, self-specialization goes a step further by incorporating domain-specific seed instructions and is further bolstered by parameter-efficient fine-tuning, as well as optional iterative refinement and retrieval components. Our goal is to guide models beyond generic alignment, directing them to generate data that are not just contextually fitting for a specialized domain but also maintain high accuracy.
We evaluate our self-specialized models within the biomedical and finance domains (20 datasets in total), and across a variety of base models that we specialize. Surprisingly, despite the simplicity of our approach, our results present a compelling case for self-specialization significantly outperforming the base models, and even larger models that are generally instruction-tuned or specifically pre-trained on the target domain. Notably, our self-specialized one based on MPT-30B (Team, 2023) for biomedicine even surpasses larger models (based on LLaMA-65B (Touvron et al., 2023a)), including the ones improved through self-alignment by leading methods (Wang et al., 2022a; Sun et al., 2023).
@@ 2 Preliminaries: Benchmarking Existing Aligned Models
To motivate our exploration of self-specialization, we first begin by addressing a fundamental question: How well do generally aligned models perform on specialized domains? While popular models, such as Alpaca (Taori et al., 2023) and Dromedary (Sun et al., 2023), have demonstrated effectiveness in following general instructions, it remains unclear whether general alignment can also elicit expertise for a certain domain.
Base Aligned Model LLaMA-65B Alpaca-65B Dromedary-65B Averaged 43.87 46.39 45.10 -Score (+2.52) (+1.23)
Investigating this, we assess the capabilities of Alpaca and Dromedary against their base model, LLaMA-65B (Touvron et al., 2023a), on a collection of benchmarks within the biomedical domain. We evaluate Alpaca as an upper bound, due to its reliance on GPT-3.5-generated datasets (Ouyang et al., 2022) via the self-instruct process (Wang et al., 2022a), unlike Dromedary, which generates instructional data from its base model. We use 10 biomedical NLP datasets (see Section 4 for details), covering a diverse set of tasks to ensure a comprehensive mix of content and also to look at the cross-task generalization, the core of instruction-tuning. Table 1 summarizes the result.
We find that both Alpaca and Dromedary have only a slight (1.2 - 2.5) advantage over LLaMA in biomedicine. While they are aligned to handle a broad set of instructions, they do not seem to effectively improve their specialized domain expertise; intuitively trading their expertise for generality given finite parameters. In light of these findings, it becomes evident that for cases where we are only interested in expert domains for all our downstream tasks, there remains a large potential for improvement beyond the generic alignment. This underscores the need for a model or approach, like self-specialization, that could potentially uncover specialization while maintaining cross-task generalizability with minimal supervision.
@@ 3 Self-Specialization
In this section, we describe our method called self-specialization illustrated in Figure 2.
3.1 Seed Demonstrations
Initially, we utilize a curated set of seed demonstrations , consisting of a triplet , comprised of instruction , a context (e.g., passage), and a response , respectively. Recognizing the difficulty of acquiring domain-specific data in real-world scenarios (Bai et al., 2021), we aim for a very minimal number of seeds: only 80 for the biomedical domain and 90 for the financial domain111While manual annotation of seed data is an assumed prerequisite for this initial step in self-alignment, we consider those numbers to be reasonable to annotate.. We leverage established datasets such as Box (Parmar et al., 2022) for seed construction to fairly ensure quality (detailed in Section 4). These seeds capture essential domain concepts but are insufficient to cover the entirety of domain knowledge. We posit that domain-relevant information, intermingled with the vast array of other information acquired during pre-training, can be effectively accessed and better utilized through our self-specialization approach, enabling these models to enhance their performance in specialized domains. Seeds provide the primary scaffold upon which subsequent domain-specific instructions are built.
@@ 3.2 Domain-Specific Instruction Generation
With the seed instructions in place, we move to generating domain-specific instructions. While these new instructions are grounded in the initial seeds, they grow to cover a comprehensive scope of the domain. Specifically, a base model , such as MPT-30B (Team, 2023) which is large enough, is prompted to produce new combinations of given a handful of seed demonstrations which are randomly sampled from the initial seeds pool. The newly formed instructions , coupled with their corresponding input contexts , shape a blueprint that the model utilizes in the following stages.
@@ 3.3 Domain-Specific Response Generation
In this phase, it is crucial for the responses not only to be correct but also to be well-aligned with the target domain. Intuitively, as this phase is conditioned on domain-specific instructions and corresponding contexts , derived from domain-specific seeds, it may be sufficient to rely on the base model itself to generate domain-specific responses. As an additional effort, we explore whether leveraging external domain-relevant knowledge would be beneficial for this case, inspired by Frisoni et al. (2022). Therefore, we optionally allow to incorporate external knowledge via a retrieval component . Specifically, forming the query as a concatenation of and , fetches top- relevant documents .
Then, each document is independently paired with the query to form a prompt to , and the final domain-specific responses are produced from the final distribution computed by marginalizing over the probabilities of each of these -combinations at each generation step.
where is a relevance score (similarity) from a retriever module and represents the language model distribution. By integrating such external information, while domain-relevant knowledge is deemed latent within LLMs, this step further encourages the generated target responses to be more nuanced and domain-specific, leading to additional improvements (Section 5.2).
@@ 3.4 Triggering Specialization
Upon establishing a set of domain-specific instructions/responses, undergoes tuning using the self-generated data, adjusting its internal parameters to cater specifically to the domain’s nuances. This step is crucial, marking the model’s transformation from being generally competent to being domain-specialized while preserving cross-task generalizability, thus resulting in the final self-aligned domain-specialized model: .
@@ 3.5 Iterative Self-Specialization
In the spirit of continuous improvement, our approach optionally supports iterative self-specialization via re-generating instructions and responses with the better-aligned model . This process has the potential of refining the model’s domain expertise with each iteration (of considering the previous iteration as base each time), iteratively improving its responses.
Biomedicine k=0 k=1 k=5 Task Dataset Base Self-Specialized Base Self-Specialized Base Self-Specialized QA BioASQ-Factoid 30.90 37.35 47.56 55.04 51.96 57.61 BioASQ-List 46.06 46.99 47.57 44.55 35.09 42.17 BioASQ-Yesno3 21.20 85.27 10.80 94.00 8.80 95.20 PubMedQA 11.98 24.16 28.89 24.87 31.69 31.31 NER AnatEM 9.63 11.99 7.57 15.76 6.59 21.25 BioNLP13CG 24.79 24.93 21.76 31.80 26.03 41.16 NCBI 18.46 14.35 27.88 43.11 17.99 46.54 RE DDI 51.00 49.40 49.20 51.60 49.38 53.40 SA Medical Drugs 35.00 65.80 11.40 54.60 11.40 32.80 DC HoC 2.44 6.01 13.91 7.61 62.84 62.65 Average 25.15 36.63 26.65 42.29 30.18 48.41
Finance k=0 k=1 k=5 Task Dataset Base Self-Specialized Base Self-Specialized Base Self-Specialized SUM EDT-Summarization 6.40 21.90 13.97 24.00 13.87 23.56 QA InsuranceQA 3.03 19.87 6.55 23.79 9.96 24.36 ConvFinQA 15.74 5.25 21.69 11.84 28.77 20.88 NER Fin3 9.94 23.93 7.53 26.95 6.80 43.87 FiNER_139 10.24 14.84 36.78 25.81 44.34 35.63 RE KPI-EDGER 11.22 31.02 43.28 53.56 49.46 63.90 SA EarningsCall 46.80 48.80 50.80 48.00 49.03 47.74 Financial_Phrasebank 23.60 73.20 9.40 47.60 29.20 68.80 FIQA-SA 44.44 56.84 58.55 61.54 61.54 70.09 CLS Gold Commodity News 21.95 43.03 61.93 55.08 38.42 61.20 Average 19.34 33.87 31.05 37.82 33.14 46.00
4 Experimental Settings
@@
For our primary evaluation, we employ various biomedical NLP datasets, most of which are curated in BigBio (Fries et al., 2022). A total of 10 different datasets are adopted to encompass a wide range of NLP tasks: Question Answering (QA), Named Entity Recognition (NER), Relation Extraction (RE), Sentiment Analysis (SA), and Document Classification (DC). Following a prior work (Parmar et al., 2022), all datasets are transformed into instructional data. Additionally, we validate our method in the financial domain to showcase its generalizability. We adopt a total of 10 diverse datasets, covering numerous NLP tasks: Summarization (SUM), QA, NER, RE, SA, and Classification (CLS), detailed in Appendix A.
@@
We employ MPT-30B (Team, 2023) as a base model for main experiments. For the retriever, we use simple yet effective BM25 (Robertson et al., 1994), assuming human-labeled data is not sufficient. For benchmarking of general-purpose aligned models, we evaluate Alpaca-65B (Taori et al., 2023) and Dromedary-65B (Sun et al., 2023) that are both based on LLaMA (Touvron et al., 2023a). In addition to MPT-30B, we adopt LLaMA-2 7B (Touvron et al., 2023b) and Falcon-40B (Almazrouei et al., 2023) to further validate the general applicability of self-specialization with different scales and base models. We additionally evaluate existing domain-specific models (Wu et al., 2023): MedLLaMA and PMC-LLaMA (Details are in Section 5.2).
@@
In our study, all tasks are approached as a unified text generation problem, aiming to assess the capabilities of generative models. In alignment with an established convention (Parmar et al., 2022), we adopt -Score as our main evaluation metric, given an early observation that Rouge-L (Lin, 2004), as shown in Table 6 in Appendix, exhibits a strong correlation with -Score.
@@
For biomedical seeds, we use data sampled from BoX (Parmar et al., 2022), encompassing 32 tasks, up to 5 instances for each dataset, resulting in a compact yet representative 80 seed samples in total, which are also used as demonstrations at inference. For optional external corpus, we leverage PubMed preprocessed in (Phan et al., 2021), which contains 30M abstracts. In the financial domain, based on our finding from biomedical experiments that showed surprising effectiveness of self-specialization relying on internal knowledge of LLMs without the external corpus, we opt not to employ an optional retrieval component to further validate the self-sufficiency of LLMs. We leverage a total of 90 seeds sampled from the 10 train sets in our corresponding benchmark datasets. We use a total of 5K synthetic data generated through our self-specialization for all experiments, unless otherwise specified. Being equipped with QLoRA (Dettmers et al., 2023) and 4-bit quantization, the model is trained using a simple Alpaca-style template (Taori et al., 2023) on a single A100, taking only a few hours for 3 epochs, resulting in a light-weight specialization module.
5 Results and Analyses
Here, we provide a set of experimental results and analyses to address relevant research questions.
5.1 Comparison with Baselines
How effective is the self-specialization of base models?